# Optimality of quasi-Monte Carlo methods and suboptimality of the sparse-grid Gauss–Hermite rule

### v1

@@ 5 Concluding remarks
These findings motivate further studies in several directions. One is to analyse sparse-grid algorithms for functions over $\mathbb{R}^{d}$ built from univariate rules other than the Gauss–Hermite rule. A notable alternative is the sparse-grid method using weighted Leja sequences [22, 48], which was proposed by Narayan and Jakeman [48]. To obtain results analogous to what we showed in this paper, it is worth noting that the fooling function we constructed in Theorems 3.1 and  3.2 depend solely on the distribution of quadrature nodes, and not on the quadrature weights. Hence, understanding the distribution of Leja points for the weight of interest would allow the construction of a corresponding fooling function.
We did not discuss dimension independence of the constant. The implied constant in our error estimate may depend on the dimension $d$, which may be large in practice. Addressing this is left for future work. A standard approach is to employ the so-called weighted spaces, as introduced by Hickernell [33], under which QMC methods can attain dimension-independent convergence rates with constants that are likewise independent of dimension, as shown by Sloan and Woźniakowski [59]. These ideas have been widely applied in the analysis of parametric partial differential equations [30, 32, 39, 40, 58]. In fact, the results in [59] show that such weights are even necessary in the setting considered there, suggesting that additional structure is required to make the constants in our bounds dimension-independent.
This paper focused on Sobolev functions. Another important regime in high dimension is that of analytic functions [8, 35, 17], where the question of optimal rates in this setting remains largely open. In establishing the lower bound for the sparse-grid Gauss–Hermite rule in this paper, a crucial component was the fooling function constructed for univariate Sobolev functions [41]. A similar technique for analytic functions is available [27, 63], which may prove useful for further analysis in this regime.

### v2

@@ 5 Concluding remarks
These findings motivate further studies in several directions. One is to analyse sparse-grid algorithms for functions over $\mathbb{R}^{d}$ built from univariate rules other than the Gauss–Hermite rule. A notable alternative is the sparse-grid method using weighted Leja sequences [undefv, undefax], which was proposed by Narayan and Jakeman [undefax]. To obtain results analogous to what we showed in this paper, it is worth noting that the fooling function we constructed in Theorems 3.1 and 3.3 depend solely on the distribution of quadrature nodes, and not on the quadrature weights. Hence, understanding the distribution of Leja points for the weight of interest would allow the construction of a corresponding fooling function.
We did not discuss dimension independence of the constant. The implied constant in our error estimate may depend on the dimension $d$, which may be large in practice. Addressing this is left for future work. A standard approach is to employ the so-called weighted spaces, as introduced by Hickernell [undefag], under which QMC methods can attain dimension-independent convergence rates with constants that are likewise independent of dimension, as shown by Sloan and Woźniakowski [undefaai]. These ideas have been widely applied in the analysis of parametric partial differential equations [undefad, undefaf, undefan, undefao, undefaah]. In fact, the results in [undefaai] show that such weights are even necessary in the setting considered there, suggesting that additional structure is required to make the constants in our bounds dimension-independent.
This paper focused on Sobolev functions. Another important regime in high dimension is that of analytic functions [undefh, undefai, undefq], where the question of optimal rates in this setting remains largely open. In establishing the lower bound for the sparse-grid Gauss–Hermite rule in this paper, a crucial component was the fooling function constructed for univariate Sobolev functions [undefap]. A similar technique for analytic functions is available [undefaa, undefaam], which may prove useful for further analysis in this regime.